At age 21 Julio begins saving ​$1 comma 200 each year until age 35 ​(15 ​payments) in an ordinary annuity paying 7.3​% annual interest compounded yearly and then leaves his money in the account until age 65​ (30 years). His friend Max begins at age 41 saving ​$2 comma 400 per year in the same type of account until age 65​ (25 payments). How much does each have in his account at age​ 65?
1
Expert's answer
2019-03-07T08:57:38-0500

For Julio the future amount is:

1) FV(15) = 1,200*((1.073^15 - 1)/0.073) = $30,860.89,

2) FV(30) = 30,860.89*1.073^30 = $255,505.41

For Max the future amount is:

FV(30) = 2,400*((1.073^25 - 1)/0.073) = $158,496.63.